OpenStudy (anonymous):

In 1 over 8 + 13 over 8 what is the shortest method of finding the GCF?

OpenStudy (anonymous):

\[\frac{1}{8}+\frac{13}{8}\] i assume you want the least common denominator, or least common multiple, not "greatest common factor"

OpenStudy (anonymous):

and since both denominators are the same, you can use 8 as your least common denominator and just add up the numerators

OpenStudy (amistre64):

\[\frac{1}{8}(1+13)\] maybe?

OpenStudy (anonymous):

Yes, I know it's 6 and mistakenly thought it was called GCF. But what's a shorter method to use for larger numbers?

OpenStudy (amistre64):

i dont think there is a short method for larger numbers. large numbers are rather abstract. computers can perform the tasks that we apply to smaller numbers on large numbers faster.

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

denominators are the same, add in the numerator \[\frac{a}{c}+\frac{b}{c}=\frac{a+b}{c}\]

OpenStudy (anonymous):

Meant to say 2. Still need a shorter method for the larger numbers though.